OpenStudy (anonymous):

Y varies jointly with P and Q. When P=3 and Q=5, Y=60. Find Y when P=2 and Q=6. A. Y=36 B. Y=48 C. Y=60 D. Y=72

OpenStudy (anonymous):

C

OpenStudy (nurali):

y=kpq 60=k(3)(5) 60=15k 4=k y=(4)(2)(6) y=-----

OpenStudy (anonymous):

I got it. Thank you very much for ur example

OpenStudy (anonymous):

Y(P+Q) since is jointly. Therefore should be 60 = (3+5)k k= 7.5

OpenStudy (anonymous):

Sorry!! Thank you for reminding me

OpenStudy (nurali):

No Problem.